OpenStudy (anonymous):

the product of two consecutive even positive integers is one hundred twenty. find the two integers

OpenStudy (anonymous):

Two consecutive integers whether even or odd is always x and x + 2 x (x + 2) = 120 Distribute x^2 + 2x = 120 One side has to equal 0 x^2 + 2x - 120 = 0 Factor (x + 12)(x - 10) = 0 Set each factor equal to 0 x + 12 = 0 or x - 10 = 0 Solve each equation x = -12 or x = 10 since it says even positive we only use the 10 Numbers are 10 and 12

OpenStudy (anonymous):

cool thanks